Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP:; phone: +1 575-835-7370


News, Travel, Visitors (McMullin)

  • Overall status: C config final days (using additional science hours in high pressure zone; working well); first science with hardware stacking; last meeting of the year next week. Reminder of EVLA operations shutdown: 24 Dec-25 Dec (32 hours); 31 Dec - 01 Jan (16 hours); C configuration will be extended through to 18 January.
  • Key dates:
  • Staff Travel:
  • Visitors:

Commissioning priorities for the week (McMullin, Chandler)

Single maintenance day week:
  • Mon-Tue, Thu-Fri:
    • 1300-1800 Testing/software development time
    • 1800-1300: Science operations
  • Wed: Maintenance day
  • Interrupt status for additional 3-bit testing; expected to be lab work only this week

Hardware status and updates (Langley, Perley, Rupen)

Software status and updates (Butler)

Commissioning Reports

Observing:Chandler: Proposed Default Frequencies by Band (Extended OSRO)

  • L-band: 2x512 MHz basebands centered at 1.25 and 1.75 GHz
  • S-band: 2x1024 MHz basebands centered at 2.5 and 3.5 GHz
  • C-band: 2x1024 MHz basebands centered at 5.0 and 6.0 GHz
  • X-band: as VLA (8396 MHz, 8524 MHz)
  • Ku-band: 2x1024 MHz basebands centered at 13.5 and 14.5 GHz
  • K-band: 2x1024 MHz basebands centered at 21.5 and 22.5 GHz
  • Ka-band: 2x1024 MHz basebands centered at 32.5 and 33.5 GHz
  • Q-band: 2x1024 MHz basebands centered at 40.5 and 41.5 GHz

Calibration: Perley: Cross-observatory flux calibrator survey


Commissioning Group updates and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!


10C-219 Irwin:

  • Short visit; returning in January; first images made; re-encountered the RFI issue (need to hanning smooth).



Demo Science Status


Misc Actions

  • [Frail/Butler]: Enable restrictions based on proximity to sun; review OST weather trending.
  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.

-- JosephMcMullin - 2010-12-06 -- JosephMcMullin - 2010-12-13
  • Ks_stacking_zoomed_marked.jpg:
    Ks stacking zoomed marked.jpg

  • Screen_shot_2010-12-14_at_9.30.43_AM.png:
    Screen shot 2010-12-14 at 9.30.43 AM.png
Topic revision: r2 - 2010-12-14, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback